OpenStudy (anonymous):

solving rational equations solve y/6=5y/3

OpenStudy (anonymous):

Since this is a fraction equal to a fraction you can cross multiply. Does that help?

OpenStudy (anonymous):

i am not sure if my answer is right it doesmt seem right

OpenStudy (anonymous):

What did you get? I'll check it.

OpenStudy (anonymous):

for some reason i got y/3=30

OpenStudy (anonymous):

When you cross multiply you get 30y=3y. Do you know where to go from there?

OpenStudy (anonymous):

i believe you divide by 3

OpenStudy (anonymous):

You can definitely do that. What do you have now?

OpenStudy (anonymous):

10y

OpenStudy (anonymous):

Good. The overall equation is 10y=y

OpenStudy (anonymous):

you can get the y's on the same side too

OpenStudy (anonymous):

yepper

OpenStudy (anonymous):

just subtract y over to the left. 10y-y=0

OpenStudy (anonymous):

help me out on the left side

OpenStudy (anonymous):

Worried I lost you. Just so you have the answer. 10y-y=0 9y=0 (now divide both sides by zero) y=0